A bag of peanuts is worth $0.20 less than a bag of cashews. Equal amounts of peanuts and cashews are used to make 48 bags of a mixture that sells for $1.15 per bag. How much is a bag of cashews worth?
Answer by Finavon(81)

Cashew C
Peanuts C-20
1 bag of each
C + C - 20 = 2*115
2C = 230+20 = 250
C = 125
Bag cashew worth $1.25
Peanuts $1.05